OpenStudy (anonymous):

A mule deer can run 1/4 of a mile in 25 seconds. At this rate, which expression can be used to determine how fast a mule deer runs in miles per hour? a. 25 seconds/ 1/4 of a mile= 36 miles per hour b. 1/4 of a mile / 25 seconds= 36 miles per hour c. 1/4 of a mile / 25 seconds= 100 miles per hour d. 25 seconds/ 1/4= 100 miles per hour @ganeshie8 @SandeepReddy

OpenStudy (fibonaccichick666):

so, what decimal of a mile is 1/4 of a mile?

OpenStudy (anonymous):

um is it 0.25

OpenStudy (fibonaccichick666):

yea, ok, so that just makes setting this up more convenient, so now you have \[\frac{.25 miles}{25 seconds}\times \frac{? seconds}{?minute}...\]

OpenStudy (daniellelovee):

how many seconds does an hour has? 60x60=360 seconds once you have that you do cross multiplication

OpenStudy (fibonaccichick666):

so now, we need to go from seconds to hours, the first step(if you don't know off the top of your head how many seconds are in an hour) is to find the conversion factor, so how many seconds are in a minute?

OpenStudy (daniellelovee):

i meant 3600 sorry

OpenStudy (anonymous):

60 seconds

OpenStudy (fibonaccichick666):

ok so now, we just fill in the question marks

OpenStudy (fibonaccichick666):

\[\frac{.25 miles}{25 seconds}\times \frac{\color\red{60} seconds}{\color\red{1}minute}...\]

OpenStudy (fibonaccichick666):

so, once we multiply that out we would be left with miles per minute as our units, can you guess how we will go from minutes to hours?

OpenStudy (anonymous):

umm.. I don't know sorry

OpenStudy (fibonaccichick666):

ok, so take a look at how we went from seconds to minutes

OpenStudy (fibonaccichick666):

we set up a fraction, that is equivalent to one. using something called a conversion factor(ie, there are 60 seconds in one minute)

OpenStudy (fibonaccichick666):

so we are trying to go from minutes to hours, can you think of what the conversion factor may be?

OpenStudy (anonymous):

60?

OpenStudy (fibonaccichick666):

60 what? you have to include units

OpenStudy (fibonaccichick666):

just use words,

OpenStudy (anonymous):

is it 60 seconds?

OpenStudy (fibonaccichick666):

well, we want to go from \(\color\orange{minutes~ to~ hours}\), so, since we have already converted seconds to minutes, we do not need another seconds.

OpenStudy (fibonaccichick666):

so what did we need to go from seconds to minutes?

OpenStudy (anonymous):

is it minutes to hours?

OpenStudy (fibonaccichick666):

please don't guess, if you don't know just say it. It's a lot harder to explain when you do.

OpenStudy (anonymous):

im sorry I didn't know :(

OpenStudy (fibonaccichick666):

it's ok, so I take it, you are a little lost on the topic? Have you been able to follow what I have said so far?

OpenStudy (anonymous):

yes so how much do we have left to complete the question

OpenStudy (fibonaccichick666):

A lot, if you do not understand, please read this http://www.purplemath.com/modules/units.htm

OpenStudy (fibonaccichick666):

it explains the process we need to perform

OpenStudy (anonymous):

I understand :)

OpenStudy (fibonaccichick666):

then what do we need to know to go from minutes to hours?

OpenStudy (anonymous):

is it like this? 60 seconds : 1 minute 60 minutes : 1 hour 1 mile : 5280 feet 1 foot : 12 inches 2.54 centimeters : 1 inch 100 centimeters : 1 meter

OpenStudy (anonymous):

@GretaKnows can you help me with this question please?

OpenStudy (fibonaccichick666):

well, those are conversion factors, but you only need to go from miles/second to miles/hr

OpenStudy (fibonaccichick666):

so you do not need all of them. This is not just one multipliction but teo

OpenStudy (anonymous):

ok :)

OpenStudy (fibonaccichick666):

so you should have enough information between the website and the work here to do the problem now. Please try

OpenStudy (anonymous):

sorry I wasn't on my computer last night!!